OpenStudy (anonymous):

Find the slope of a line parallel to the line through the given points. P(-2, 4), Q(6, -2)

OpenStudy (anonymous):

-3/4 3/4 4/3

OpenStudy (jim766):

find the slope y2 -y1/ x2-x1 -2 - 4/ 6 --2

OpenStudy (jim766):

parallel lines have the same slopes...so the slope of that line will be the slope of the parallel line

OpenStudy (anonymous):

-3/4?

OpenStudy (jim766):

-2-4 = -6 6--2 = 8 minus a negative is same as addition -6/8 = -3/4

OpenStudy (jim766):

makes sense?

OpenStudy (anonymous):

yes, thank you!